### v1

@@ 1.2 Contribution
To determine the optimal weights matrix $\hat{W}$ in phase (1) of the procedure, the network’s activation function must be invertible and take values in the range $[-1,1]$. Therefore, in the examples presented in this paper, we use the Erf activation function. The key distinction between our ENR-ELM and the traditional ELM lies in the choice of the hidden layer weight matrix, which is data-dependent rather than random in our approach. Although the spectral decomposition of the Gram matrix is computationally intensive, this cost is offset by the linear complexity of the second phase, making the overall procedure more efficient compared to the traditional ELM, particularly in terms of model selection.
@@ 2.2 ELM
where $W\in\mathbb{R}^{n\times n_{0}}$ is a random matrix with independent entries $W_{ij}\sim{\cal{N}}(0,1/\sqrt{n_{0}})$, and $\sigma:\mathbb{R}\to\mathbb{R}$ is referred to as the activation function, applied entry-wise to the vector $W\,\mathbb{x}$. Given the activation function $\sigma$, the weights matrix $W$, and a weights vector $\boldsymbol{\beta}\in\mathbb{R}^{n}$, a shallow fully connected neural network is the following function
@@ 2.3 Neural Network Gaussian Process Kernel
Consider a stochastic fully connected L-layered neural network, as the one given in eq. (7) where all elements of $\mathcal{W}$ are independent Gaussian variables with zero mean and variance $1/n$, except in the first hidden layer where the variance is $1/n_{0}$, then the NNGP kernel is the covariance of the limiting Gaussian process defined as
@@ 3 The proposed method
Since the columns of $U$ are by construction orthonormal, their entries are within the range $[-1,1]$; assuming that the activation function is invertible and takes values in $[-1,1]$, we obtain the solution of equation (9) in a closed form, i.e. $\hat{W}=\sigma^{-1}(U^{\top})\,X^{\top}\,(XX^{\top})^{-1}$, where $\sigma^{-1}$ is intended to be applied entry-wise. After evaluating the optimal matrix $\hat{W}$ we have two approaches. In the first approach, which we call A-ENR-ELM (approximated ENR-ELM), we learn the output layer coefficients using the OLS to regress $y$ on the matrix $U$, which is a proxy for the matrix $\sigma(\hat{W}X)^{\top}$. This approach computes the final layer coefficients $\hat{\boldsymbol{\beta}}=U^{\top}y$ with linear computational cost. In the second approach, we learn the output layer coefficients using the classical incremental forward stagewise regression of $y$ on matrix $S=\sigma(\hat{W}X)^{\top}$, for that reason we call this second approach I-ENR-ELM (incremental ENR-ELM). In this second approach, each neuron (which corresponds to a column of matrix $S$) is partially added to the estimator incrementally until some convergence criterion is achieved.
@@ 4.2 Results on synthetic datasets
Each synthetic dataset was created by generating $T\in\{300,1200\}$ pairs $\{(\mathbb{x}_{l},y_{l})\}_{l=1,...,T}$ where $\mathbb{x}_{l}\in\mathbb{R}^{n_{0}}$ with dimension $n_{0}=20$ or $n_{0}=80$ and $y_{l}=f(\mathbb{x}_{l})+\epsilon_{l}$, with $f(\cdot)$ being a given function and $\epsilon_{l}$ are independent and identically distributed Gaussian variables with zero mean and fixed variance $\sigma^{2}$. The generic input data vectors $\mathbb{x}$ were independently sampled from the following three different types of distribution: Uniform, where the vector components are independent and uniformly distributed in the interval $[-2\pi,2\pi]$; Indep. Gaussian, where the vector components are independent and identically distributed as a standard Gaussian variable; Toepl. Gaussian, where the vector $\mathbb{x}$ is distributed as $n_{0}$-variate Gaussian variable with zero mean and covariance given by $cov(x_{i},x_{j})=\rho^{|i-j|}$ with $\rho=0.8$. We consider two types of function $f(\cdot)$: Linear, where. $f(\mathbb{x})=\mathbb{\alpha}^{\top}\mathbb{x}+\beta$ where the components of $\mathbb{\alpha}$ and $\beta$ are independent and uniformly distributed in the interval $[-2,2]$, and Shallow NN, where $f(\mathbb{x})$ is the output obtained by a Neural Network with one hidden layer composed of 100 neurons whose hidden layer weights are drawn from a Gaussian distribution with zero mean and variance $1/n_{0}$, while the output layer weights are generated according to a Gaussian distribution with zero mean and variance $1/100$. Finally, we fix the variance of the noise $\sigma^{2}$ such that the Signal to Noise Ratio (SNR), defined as the ratio between the variance of the unknown signal and the variance of the noise, results to be $SNR=2$ or $SNR=10$, the first case representing a strong noise case, the second case representing a weak noise case.

### v2

@@ Abstract
This preprint has not undergone peer review or any post‑submission improvements or corrections. The Version of Record of this article is published in Neural Computing and Applications, and is available online at https://doi.org/10.1007/s00521-025-11519-5
@@ 1.2 Contribution
To determine the optimal weights matrix $\hat{W}$ in phase (1) of the procedure, the network’s activation function must be invertible and take values in the range [1,1][-1,1][ - 1 , 1 ]. Therefore, in the examples presented in this paper, we use the Erf activation function. The key distinction between our ENR-ELM and the traditional ELM lies in the choice of the hidden layer weight matrix, which is data-dependent rather than random in our approach. Although the spectral decomposition of the Gram matrix is computationally intensive, this cost is offset by the linear complexity of the second phase, making the overall procedure more efficient compared to the traditional ELM, particularly in terms of model selection.
@@ 2.2 ELM
where $W\in\mathbb{R}^{n\times n_{0}}$ is a random matrix with independent entries $W_{ij}\sim{\cal{N}}(0,1/n_{0})$, and $\sigma:\mathbb{R}\to\mathbb{R}$ is referred to as the activation function, applied entry-wise to the vector $W\,\mathbb{x}$. Given the activation function $\sigma$, the weights matrix $W$, and a weights vector $\boldsymbol{\beta}\in\mathbb{R}^{n}$, a shallow fully connected neural network is the following function
@@ 2.3 Neural Network Gaussian Process Kernel
[…]
@@ 3 The proposed method
Since the columns of $U$ are by construction orthonormal, their entries are within the range [1,1][-1,1][ - 1 , 1 ]; assuming that the activation function is invertible and takes values in [1,1][-1,1][ - 1 , 1 ], we obtain the solution of equation (9) in a closed form, i.e. $\hat{W}=\sigma^{-1}(U^{\top})\,X^{\top}\,(XX^{\top})^{-1}$, where $\sigma^{-1}$ is intended to be applied entry-wise. After evaluating the optimal matrix $\hat{W}$ we have two approaches. In the first approach, which we call A-ENR-ELM (approximated ENR-ELM), we learn the output layer coefficients using the OLS to regress $y$ on the matrix $U$, which is a proxy for the matrix $\sigma(\hat{W}X)^{\top}$. This approach computes the final layer coefficients $\hat{\boldsymbol{\beta}}=U^{\top}y$ with linear computational cost. In the second approach, we learn the output layer coefficients using the classical incremental forward stagewise regression of $y$ on matrix $S=\sigma(\hat{W}X)^{\top}$, for that reason we call this second approach I-ENR-ELM (incremental ENR-ELM). In this second approach, each neuron (which corresponds to a column of matrix $S$) is partially added to the estimator incrementally until some convergence criterion is achieved.
@@ 4.2 Results on synthetic datasets
Each synthetic dataset was created by generating $T\in\{300,1200\}$ pairs $\{(\mathbb{x}_{l},y_{l})\}_{l=1,...,T}$ where $\mathbb{x}_{l}\in\mathbb{R}^{n_{0}}$ with dimension $n_{0}=20$ or $n_{0}=80$ and $y_{l}=f(\mathbb{x}_{l})+\epsilon_{l}$, with $f(\cdot)$ being a given function and $\epsilon_{l}$ are independent and identically distributed Gaussian variables with zero mean and fixed variance $\sigma^{2}$. The generic input data vectors $\mathbb{x}$ were independently sampled from the following three different types of distribution: Uniform, where the vector components are independent and uniformly distributed in the interval $[-2\pi,2\pi]$; Indep. Gaussian, where the vector components are independent and identically distributed as a standard Gaussian variable; Toepl. Gaussian, where the vector $\mathbb{x}$ is distributed as $n_{0}$-variate Gaussian variable with zero mean and covariance given by $cov(x_{i},x_{j})=\rho^{|i-j|}$ with $\rho=0.8$. We consider two types of function $f(\cdot)$: Linear, where. $f(\mathbb{x})=\mathbb{\alpha}^{\top}\mathbb{x}+\beta$ where the components of $\mathbb{\alpha}$ and $\beta$ are independent and uniformly distributed in the interval [2,2][-2,2][ - 2 , 2 ], and Shallow NN, where $f(\mathbb{x})$ is the output obtained by a Neural Network with one hidden layer composed of 100 neurons whose hidden layer weights are drawn from a Gaussian distribution with zero mean and variance $1/n_{0}$, while the output layer weights are generated according to a Gaussian distribution with zero mean and variance 1/1001/1001 / 100. Finally, we fix the variance of the noise $\sigma^{2}$ such that the Signal to Noise Ratio (SNR), defined as the ratio between the variance of the unknown signal and the variance of the noise, results to be $SNR=2$ or $SNR=10$, the first case representing a strong noise case, the second case representing a weak noise case.